I have two equations. The answers for both X & Y are the same. I have tried to figure it out and have been unable to get the answer. Here are the equations.
2x + 3y = 33
x + 5y = 27

Answer by LilSkittleMd(119)

2x + 3y = 33
x + 5y = 27
Solve the second equation for x
x=27-5y
Substitute the value of x into the first equation
2(27-5y)+3y=33
Distribute
54-10y+3y=33
Combine like terms
54-7y=33
Subtract 54 from both sides
-7y=-21
Divide each side by -7
y=3
Plug in the value of y into one of the original equations
x+5(3)=27
x+15=27
x=12
x=12 and y=3